OpenStudy (anonymous):

Find the x-coordinates of any relative extrema and inflection point(s) for the function f(x) = 6x(1/3) + 3x(4/3). You must justify your answer using an analysis of f ′(x) and f ′′(x).

OpenStudy (freckles):

have you found f' and f'' yet?

OpenStudy (freckles):

also do you mean f(x)=6x^(1/3)+3x^(4/3)?

OpenStudy (freckles):

if you have found f' and f'' please share them

OpenStudy (mathmale):

Taking calculus, as you apparently are, you should develop skills in presenting problems correctly. Freckles, above, has tried to find out whether you mean exponentiation. One way of presenting math'l equations properly is through Equation Editor. Your \[f(x)=6x^(1/3)+3x^(4/3)\]

OpenStudy (mathmale):

becomes \[ f(x)=6x ^{1/3}+3x ^{4/3}\]

OpenStudy (mathmale):

Please differentiate this to obtain the first and second derivatives of f(x).

OpenStudy (anonymous):

ok an dyes thats what i meant

OpenStudy (freckles):

power rule is your friend here You will be using (x^n)'=n*x^(n-1) like crazy here

OpenStudy (anonymous):

f'x = (4x+2)/x^2/3

OpenStudy (freckles):

that looks good for f'

OpenStudy (anonymous):

f"(x) = (4x-4)/3x^5/3

OpenStudy (freckles):

that looks great too

OpenStudy (freckles):

the domain for our given function is all real numbers now to find critical numbers you need to find when f'=0 and when f' does not exist

OpenStudy (freckles):

we will need to do this to find the relative extrema

OpenStudy (freckles):

so when f'=0 and when f' does not exist we have the following two equations to solve: \[4x+2=0 \\ x^{\frac{2}{3}}=0\] You need to solve both of these equations (we will do the same thing for f'' to find the possible inflection points )

OpenStudy (anonymous):

th first one is x=2/4

OpenStudy (freckles):

well 4x+2=0 4x=-2 x=-2/4 or reducing gives x=-1/2 --- last equation should be easy x^(2/3) is 0 when x is 0 draw number line:  I prefer to use the first derivative to see when the function is increasing and decreasing to determine the max and min we will test the 3 intervals above by choosing a number representative for each interval

OpenStudy (freckles):